OpenStudy (anonymous):

–x + 3y = 9 than please describe how to graph using slope intercept form

OpenStudy (saifoo.khan):

When x intercept put x = 0 when y intercept put y = 0

OpenStudy (anonymous):

nd then?

OpenStudy (saifoo.khan):

try solving..

OpenStudy (anonymous):

for?

OpenStudy (anonymous):

y=x/3+3

OpenStudy (saifoo.khan):

When x intercept put x = 0 -x + 3y = 9 -x + 3(0) = 9 solve. when y intercept put y = 0 This one as well.

OpenStudy (anonymous):

noo... thats not what its asking.. thanks anyway!

OpenStudy (saifoo.khan):

that's the thing how u are going to graph it.

hero (hero):

–x + 3y = 9 in slope intercept form is \[y = \frac{9+x}{3}\]

hero (hero):

where the slope is 1/3 and the intercept is 3

OpenStudy (anonymous):

this is waht i got... y=mx+b 3y=x+9 y=x/3+3

OpenStudy (saifoo.khan):

\[y = \frac93 + \frac x3\] \[y = \frac x3+3\]

hero (hero):

lol, why'd it take you so long saifoo?

OpenStudy (anonymous):

hero click on waht i just posted

OpenStudy (saifoo.khan):

i was in another tab. :p